# Enhancing Reinforcement Learning in Sensor Fusion: A Comparative Analysis of Cubature and Sampling-based Integration Methods for Rover Search Planning

### v1

@@ I Introduction
A lot of research effort is dedicated to increasing the robustness of ML algorithms against noise [6], or even creating new ones specifically for noisy training [7]. However, in the realm of reinforcement learning noise can be beneficial. Action noise, which is the addition of randomness to the action selected by the algorithm, can be desirable to encourage exploration of the action space. Using Gaussian or Ornstein-Uhlenbeck noise has been shown to significantly improve training across popular reinforcement algorithms such as Soft Actor-Critic [8] or Deep Deterministic Policy Gradient [9] when coupled with scheduled noise reduction [10]. However, this scheduling shows that noise must be minimized to fine-tune the model. In comparison, observation noise in reinforcement learning, akin to poor-quality labelled data for supervised learning, is known to make training unstable [11]. Noise in ML, when not controllable, is undesirable as it increases the difficulty in successfully learning a model.
@@
where $\vec{x}\in\mathbb{R}^{2}$, $f$ is the integrand, and $H$ is the collection of $s$-simplices. In search path planning, the integrand is commonly a Probability Distribution Map (PDM), and $H$ is the path buffered by the sensor footprint $r$. A widely used method of integration is a sampling-based approach which calculates the integration result through grid sampling of the PDM and then calculating the summation of pixels within $H$. A different approach uses cubature [12] to achieve a similar result. This relies on the transformation of $H$ into a sum of computable simplices such as triangles or rectangles. The works by [13] implement methods for 15 primitive regions that any simplex can be dissected into, like the aforementioned triangle and rectangle. The integration result is then the sum of cubature results for each simplex.
@@
Search planning typically is applied in situations where exhaustive search (also known as coverage planning) is infeasible due to mission constraints such as limited resources. One such situation where a reduced vehicle lifespan is the limiting factor is searching the Martian surface for items of scientific interest based on a PDM. In 2020, NASA’s Mars Mission utilized two robots: the Perseverance rover and the Ingenuity helicopter - opening the door to the use of collaborative robotic teams for planetary exploration [14]. The use of multi-rover teams for planetary exploration missions has been considered in terms of both surface [15] and cave [16] exploration. By using a team of planetary exploration rovers, the data collection capabilities of a mission can be increased due to the extended sensor footprint.
@@ II-A Mission Planning
The PDM is modelled as a sum of $G$ bivariate Gaussians[17] such that a point on the surface of Mars at coordinate $\vec{x}\in\mathbb{R}^{2}$ has a probability of containing the meteor p
@@ II-C2 Cubature Integration
The simplex $H$ must first be subdivided into a set of the 15 primitive simplices [13], with the simplest and most ubiquitous of these being the triangle. As $H$ is likely to have many holes and unlikely to be convex, the constrained Delaunay triangulation [chew_constrained_1989] is used to get the set of triangles $T$.
@@ III Results
Figure 6: The relative error as $N$ increases showing a trend towards an exponential increase as $N\rightarrow\infty$.
@@ IV Conclusion
Overall, the results of this study highlight the superiority of the cubature approach in terms of accuracy and computational efficiency, particularly in scenarios requiring coordinated search missions like those encountered in reinforcement learning and embedded systems for autonomous navigation. However, it’s important to acknowledge the limitations of this work. While the focus was on comparing two prevalent integration methods, numerous other techniques warrant exploration. Additionally, the requirement of a continuous integrand for the cubature method may pose challenges in certain scenarios, though potential solutions using approximations have been discussed. By recognizing these limitations, future research can build upon these findings to further enhance integration techniques for a wide range of applications.

### v3

@@ I Introduction
While significant progress has been made in developing robust ML algorithms [6, 7], the impact of noise in reinforcement learning remains a critical challenge [8]. In the realm of reinforcement learning noise can be beneficial. Action noise, which is the addition of randomness to the action selected by the algorithm, can be desirable to encourage exploration of the action space. Using Gaussian or Ornstein-Uhlenbeck noise has been shown to significantly improve training across popular reinforcement algorithms such as Soft Actor-Critic [9] or Deep Deterministic Policy Gradient [10] when coupled with scheduled noise reduction [11]. However, this scheduling shows that noise must be minimized to fine-tune the model. In comparison, observation noise in reinforcement learning, akin to poor-quality labelled data for supervised learning, is known to make training unstable [8]. Noise in the reward channel for reinforcement learning is particularly undesirable as it can lead to actions being misinterpreted. [12] highlights multiple scenarios in which the agent may take advantage of a reward function leading to reward corruption and a stagnation in learning on a local maxima. Whilst a solution to this was proposed in [12], a much easier approach is to minimize the chances of this happening in the first place by implementing a robust method wth as little noise as possible. Overall noise in ML, when not controllable, is undesirable as it increases the difficulty in successfully learning the policy.
@@
where $\vec{x}\in\mathbb{R}^{2}$, $f$ is the integrand, and $H$ is the collection of $s$-simplices. In search path planning, the integrand is commonly a Probability Distribution Map (PDM), and $H$ is the path buffered by the sensor footprint $r$. A widely used method of integration is a sampling-based approach which calculates the integration result through grid sampling of the PDM and then calculating the summation of pixels within $H$. A different approach uses cubature [13] to achieve a similar result. This relies on the transformation of $H$ into a sum of computable simplices such as triangles or rectangles. cubpack++ by [14], implement methods for 15 primitive regions that any simplex can be dissected into, like the aforementioned triangle and rectangle. The integration result is then the sum of cubature results for each simplex. Another approach by [15], r2d2lri, uses a sequence of embedded lattice rules and claims to result in little computational overhead. Whilst [16] found r2d2lri to be faster during integration over a infinite domain, cubpack++ proved to be more adaptable to different scenarios without a massive speed penalty. For the rest of this work, cubpack++ will be used for for its adaptability and adaptive properties.
@@
Search planning typically is applied in situations where exhaustive search (also known as coverage planning) is infeasible due to mission constraints such as limited resources. One such situation where a reduced vehicle lifespan is the limiting factor is searching the Martian surface for items of scientific interest based on a PDM. In 2020, NASA’s Mars Mission utilized two robots: the Perseverance rover and the Ingenuity helicopter - opening the door to the use of collaborative robotic teams for planetary exploration [17]. The use of multi-rover teams for planetary exploration missions has been considered in terms of both surface [18] and cave [19] exploration. By using a team of planetary exploration rovers, the data collection capabilities of a mission can be increased due to the extended sensor footprint. Whilst this study was focused on exploratory Mars rovers, search planning has applications within other domains such as in wilderness [20] or marine [21] search and rescue.
@@ II-A Mission Planning
The PDM is modelled as a sum of $G$ bivariate Gaussian [22] such that a point on the surface of Mars at coordinate $\vec{x}\in\mathbb{R}^{2}$ has a probability of containing the meteor p
@@ II-C2 Cubature Integration
In order to convert the simplex $H$ into something usable for cubpack++, it must first be subdivided into a set of the 15 primitive simplices, with the simplest and most ubiquitous of these being the triangle. As $H$ is likely to have many holes and unlikely to be convex, the constrained Delaunay triangulation [36] is used to get the set of triangles $T$.
@@ III Results
Figure 6: The relative time as $N$ increases showing a trend towards an exponential increase as $N\rightarrow\infty$.
@@ IV Conclusion
Overall, the results of this study highlight the superiority of the cubature approach in terms of accuracy and computational efficiency, particularly in scenarios requiring coordinated search missions like those encountered in reinforcement learning and embedded systems for autonomous navigation. However, it’s important to acknowledge the limitations of this work. While the focus was on comparing two prevalent integration methods, numerous other techniques warrant exploration. As highlighted in Sect. I, other potentially faster but less adaptable cubature algorithms exist which could result in an even great performance increase. Additionally, the requirement of a continuous integrand for the cubature method may pose challenges in certain scenarios, though potential solutions using approximations have been discussed. Furthermore, it is important to note that the influence of problem characteristics (deterministic or stochastic, complexity, etc.) on the comparison remains unexplored. By recognizing these limitations, future research can build upon these findings to further enhance integration techniques for a wide range of applications.